OpenStudy (anonymous):

an athletic has 246 members of these 164 lift weights and 208 perform cardiovascular exercises regularly all members do at least one of these activities. How many do both?

OpenStudy (anonymous):

Any body know this

hartnn (hartnn):

let A=lift weights,B=perform cardiovascular exercises from the data ,n(A)=164,n(B)=208,n(AUB)=246,to find n(A and B) the formula is n(A and B)=n(A)+n(B)-n(AUB) so n(A and B)=164+208-246 can u solve further??

OpenStudy (anonymous):

yes would it be 126 but that is not one of my choices

OpenStudy (anonymous):

then your text is wrong, or you relayed the information wrong.